OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (0, 1) and a directrix of y = -1

OpenStudy (anonymous):

I think the answer is f(x)=1/4x^2 but i am not sure.

OpenStudy (anonymous):

no one likes these conic sections question do they?

OpenStudy (mertsj):

If the focus is (0,1) and the directrix is y=-1, then the vertex is (0,0) and the parabola is concave upward. So it's equation is: y = ax^2

OpenStudy (anonymous):

btw your answer is right, since the vertex is at \((0,0)\) which in one unit away from the directrix, you have \(4py=x^2\) where \(p=1\) so you can write it as \[y=\frac{1}{4}x^2\]

OpenStudy (anonymous):

Thank you! So in that case, if the focus is (-2,4) and the directrix is y= 6 then the vertex =(-2,5) and the parabola is concave downward, so its equation would be \[f(x)=1/4(x+2)^2+5\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

I'm sorry, the equation would be: \[f(x)=-1/4(x+2)^2+5\]

OpenStudy (mertsj):

yes